OpenStudy (anonymous):

a student gets 20 out of 25 questions on a test correct. what grade does the student make?

OpenStudy (anonymous):

20 / 25 = x / 100 cross multiply 25x = 2000 x = 2000/25 x = 80

OpenStudy (amistre64):

20:25 *4 80:100